OpenStudy (anonymous):

A large asteroid crashed into a moon of a planet, causing several boulders from the moon to be propelled into space toward the planet. Astronomers were able to measure the speed of one of the projectiles. The distance (in feet) that the projectile traveled each second, starting with the first second, was given by the arithmetic sequence 26, 44, 62, 80,.... Find the total distance that the projectile traveled in seven seconds. A) 534 B) 560 C) 212 D) 426

OpenStudy (anonymous):

DON'T JUST CHOOSE SOME ANSWER EXPLAIN IT!

OpenStudy (anonymous):

@campbell_st try this please

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow a_{n} = a_{1} + (n - 1)d }\)

Parth (parthkohli):

Common difference is d(18) a1 is the first term: 26

Parth (parthkohli):

You have to find a7

OpenStudy (campbell_st):

ok... you know the 1st term a = 26 and you are asked to find the sum of the 1st 7 terms this needs the formula \[s _{n} = \frac{n}{2} [2a + (n-1)d]\] the common difference is \[T _{2} - T _{1} =\] check its the same as \[T _{3}-T _{2}=\] then n = 7, a = 26 and your d value substitute and evaluate

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow a_{7} = 26 + 18(7 - 1) }\)

Parth (parthkohli):

So, a7 = 26 + 18(6) a7 = 26 + 108 a7 = 134

Parth (parthkohli):

Got it?

Parth (parthkohli):

The common difference in an arithmetic sequence is: \(\Large \color{purple}{\rightarrow a_{n+ 1} - a_{n} }\)

OpenStudy (anonymous):

134 isnt any of those answers and those r the only answers they give me on my optional study guide

Parth (parthkohli):

Oops

Parth (parthkohli):

@campbell_st Help please, I'm bad at sequences

OpenStudy (campbell_st):

there is an altenative formula for sum of an arithmetic series \[s _{n} = \frac{n}{2}[a + l]\] this is a = 1st term l = (last term term 7) which is what parth has found..

OpenStudy (campbell_st):

so to use parth's informtion a = 26 l = 134 \[s _{7} = \frac{7}{2}(26 + 134)\]

OpenStudy (campbell_st):

either formula will give the same answer

Parth (parthkohli):

So was I correct?

OpenStudy (campbell_st):

no parth... all you have done is found the 7th term... the question is asking for the sum of the 1st 7 terms...

OpenStudy (campbell_st):

my answer was 560

Parth (parthkohli):

Oh yes, it said series.

OpenStudy (campbell_st):

you were part way to a solution

Parth (parthkohli):

I didn't read the "TOTAL DISTANCE"

OpenStudy (campbell_st):

hope this helps algebra...